### v2

@@ Abstract
In this paper, we study the risk sharing problem among multiple agents using Lambda Value-at-Risk as their preference functional, under heterogeneous beliefs, where beliefs are represented by several probability measures. We obtain semi-explicit formulas for the inf-convolution of multiple Lambda Value-at-Risk measures under heterogeneous beliefs and the explicit forms of the corresponding optimal allocations. To show the impact of belief heterogeneity, we consider three cases: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs. For those cases, we find more explicit expressions for the inf-convolution, showing the influence of the relation of the beliefs on the inf-convolution. Moreover, we consider, in a two-agent setting, the inf-convolution of one Lambda Value-at-Risk and a general risk measure, including expected utility, distortion risk measures and Lambda Value-at-Risk as special cases, with differing beliefs. The expression of the inf-convolution and the form of the optimal allocation are obtained. In all above cases we demonstrate that trivial outcomes arise when both belief inconsistency and risk tolerance are high. Finally, we discuss risk sharing for an alternative definition of Lambda Value-at-Risk.
@@ 1 Introduction
The relation between the beliefs plays a crucial role in the risk sharing problem; see e.g., Chi (2019). To study the impact of belief heterogeneity, we consider three cases in Section 3: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs. For homogeneous beliefs, we show that the inf-convolution of $\Lambda\mathrm{VaR}$ is still a $\Lambda\mathrm{VaR}$ for general $\Lambda$ functions, extending the results in Liu (2024) and Xia and Hu (2024), where the monotonicity of the $\Lambda$ functions is typically required. In particular, our results include the case that some $\Lambda\mathrm{VaR}$ have increasing $\Lambda$ functions and others have decreasing $\Lambda$ functions, demonstrating that the agents may have relatively different risk appetites. The conditional beliefs reflect the asymmetric information obtained by the agents such that each agent has slightly different concern. Under this setup, we also obtain an explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$, showing that the inf-convolution of $\Lambda\mathrm{VaR}$ under conditional beliefs is also a $\Lambda\mathrm{VaR}$ under a new conditional belief. Conditional beliefs are closely related to the conditional risk measures proposed in the literature to measure systemic risk, such as $\mathrm{CoVaR}$ (Adrian and Brunnermeier (2016) and Girardi and Tolga Ergün (2013)) and $\mathrm{CoES}$ (Mainik and Schaanning (2014)). The risk sharing problem in this case can be interpreted as risk sharing with some Co-risk measures ($\mathrm{Co\Lambda\mathrm{VaR}}$, a new conditional risk measure proposed in Section 3). For the third case, we consider the risk sharing between two agents and suppose one probability measure is absolutely continuous with another one. Under this assumption, we find a more explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$. Note that we also give the results for the inf-convolution of $\mathrm{VaR}$ with conditional beliefs and absolutely continuous beliefs respectively, which are also new to the literature.
In Section 4, we study the inf-convolution of one $\Lambda\mathrm{VaR}$ and a general monotone risk measure with belief heterogeneity. The expression of the inf-convolution and the form of the optimal allocation are derived. Then we consider two cases: conditional beliefs and absolutely continuous beliefs. For conditional beliefs, we obtain explicit expressions for the inf-convolution of one $\Lambda\mathrm{VaR}$ and one distortion risk measure/expected utility/$\Lambda\mathrm{VaR}^{+}$. For absolutely continuous beliefs, we find the formula for the inf-convolution of one $\Lambda\mathrm{VaR}$ and one $\mathrm{ES}$/expected utility. In Section 5, we consider the inf-convolution of one $\Lambda\mathrm{VaR}^{+}$ (an alternative definition of Lambda Value-at-Risk) and a general monotone risk measure with belief heterogeneity. We obtain the expression of the inf-convolution, which is complicated, and the forms of the optimal allocation. Our results include the inf-convolution of two $\Lambda\mathrm{VaR}^{+}$ under heterogeneous beliefs as a special case. The results here are not explicit and very different from the previous sections, as they involve the shape of the $\Lambda$ function and the best case of risk aggregation under dependence uncertainty. This means that $\Lambda\mathrm{VaR}^{+}$ is very different from $\Lambda\mathrm{VaR}$ for its application within the risk sharing problem.
@@
The notation and key definitions are given in Section 2 and all the proofs are delegated to Appendix A-D.
@@ 2 Notation and Definitions
For a given atomless probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $L^{\infty}$ denote the collection of all bounded random variables. Moreover, let $\mathcal{X}$ be a set of random variables containing $L^{\infty}$. We suppose $\mathcal{X}$ has good enough properties to conduct our study, such as $\mathcal{X}=L^{p}$ for some $p\geqslant 0$, where $L^{p}$ represents the set of all random variables with finite $p$-th moments. Throughout this paper, we assume that the probability measures $\mathbb{Q},\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ live on $(\Omega,\mathcal{F})$ and the corresponding probability spaces are atomless. For any $X\in\mathcal{X}$, a positive value of $X$ represents a financial loss. The distribution function of $X$ under $\mathbb{Q}$ is denoted as $F_{X}^{\mathbb{Q}}$. For a mapping $\rho:\mathcal{X}\to\mathbb{R}$, we say $\rho$ is law-invariant under $\mathbb{Q}$ if for all $X,Y\in\mathcal{X}$,
@@
where $\overset{\mathbb{Q}}{=}$ stands for equality in distribution under $\mathbb{Q}$; $\rho$ is monotone if $X\leqslant Y$ implies $\rho(X)\leqslant\rho(Y)$; and $\rho$ is cash-additive if $\rho(X+c)=\rho(X)+c$ for $X\in\mathcal{X}$ and $c\in\mathbb{R}$. We say $\rho$ is a monetary risk measure if it is monotone and cash-additive. For more details on risk measures, one can refer to Chapter 4 of Föllmer and Schied (2016). For simplicity, throughout the paper, we use the notation $\rho^{\mathbb{Q}}$ to represent that $\rho$ is law-invariant under $\mathbb{Q}$. For $X\in\mathcal{X}$, $F_{X}^{\mathbb{Q},-1}$ represents its left-quantile under $\mathbb{Q}$, which is defined by
@@
with the convention that $\inf\emptyset=\infty$. For any $X\in\mathcal{X}$, we denote by $U_{X}^{\mathbb{Q}}$ a uniform random variable on $[0,1]$ under $\mathbb{Q}$ such that $X=F_{X}^{\mathbb{Q},-1}(U_{X}^{\mathbb{Q}})$ a.s. under $\mathbb{Q}$. The existence of such $U_{X}^{\mathbb{Q}}$ for any random variable $X$ is guaranteed by e.g., Lemma A.32 of Föllmer and Schied (2016).
@@
In this paper, we suppose the agents may have heterogeneous beliefs, represented by some probability measures $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$. The inf-convolution of the risk measures $\rho_{1}^{\mathbb{Q}_{1}},\dots,\rho_{n}^{\mathbb{Q}_{n}}$ is the mapping $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}:\mathcal{X}\to[-\infty,\infty)$, defined as
@@
Note that $\rho_{i}$ is law-invariant under $\mathbb{Q}_{i}$, where $\mathbb{Q}_{i}$ represents the belief of the i-th agent.
An $n$-tuple $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is called an optimal allocation of $X$ for $(\rho_{1}^{\mathbb{Q}_{1}},\dots,\rho_{n}^{\mathbb{Q}_{n}})$ if $\sum_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X_{i})=\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)$. The inf-convolution $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}(X)$ can be interpreted as the smallest possible aggregate capital for the total risk $X$ in financial system, if $\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ represents the capital charge for the i-th financial institution to hold the risky position $X_{i}$. More economic interpretations on the inf-convolution can be found in e.g., Delbaen (2012), Rüschendorf (2013) and Embrechts et al. (2018).
For the risk measures $(\rho_{1},\dots,\rho_{n})$ and a total risk $X$, an allocation $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is Pareto-optimal if for any other allocation $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$, $\rho_{i}(Y_{i})\leqslant\rho_{i}(X_{i})$ for all $i=1,\dots,n$ implies $\rho_{i}(Y_{i})=\rho_{i}(X_{i})$ for all $i=1,\dots,n$. For finite-valued monetary risk measures, it is shown in Embrechts et al. (2018) that an allocation is optimal if and only if it is Pareto-optimal. Note that $\Lambda\mathrm{VaR}$ (defined in (2)) does not satisfy cash-additivity; see Frittelli et al. (2014) and Bellini and Peri (2022). Hence it is not a monetary risk measure. However, one can still show that the optimal allocation of the inf-convolution of $\Lambda\mathrm{VaR}$ is Pareto-optimal.
@@
where $\inf\emptyset=\infty$ and $\sup\emptyset=-\infty$. Note that $\Lambda\mathrm{VaR}(X)=\Lambda\mathrm{VaR}^{+}(X)$ if $\Lambda$ is increasing; Otherwise, it may not be true; see Proposition 6 of Bellini and Peri (2022). We refer to Bellini and Peri (2022) for two other definitions of $\Lambda\mathrm{VaR}$. If $\Lambda$ is a constant, then $\Lambda\mathrm{VaR}^{\mathbb{Q}}$ boils down to $\mathrm{VaR}^{\mathbb{Q}}$, i.e., $\mathrm{VaR}^{\mathbb{Q}}$ at level $p\in[0,1)$ is given by
@@
The conclusion in Proposition 1 implies that for any two monetary risk measures $\rho_{1}$ and $\rho_{2}$, $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)=-\infty$ if $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ are mutually singular. This means that if the beliefs of agents are too divergent, then each one of them can get unreasonably good deal by transferring risks between them. Then, the fact that the inf-convolution is $\infty$- ∞ reflects that agents can always make their position better and hence there is no optimal allocation. In contrast, if $\mathbb{Q}_{1}=\mathbb{Q}_{2}$, then $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)>-\infty$ for many monetary risk measures $\rho_{1}$ and $\rho_{2}$; see e.g., Embrechts et al. (2018) for quantile-based risk measures and Filipović and Svindland (2008) for convex risk measures. Note that the conclusion in Proposition 1 can be easily extended to the case with $n$ agents for $n\geqslant 3$.
@@ 3 Inf-convolution of multiple $\Lambda\mathrm{VaR}$
In this section, we investigate the inf-convolution of $n$ Lambda Value-at-Risk with heterogeneous beliefs represented by $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$, respectively. Suppose $\mathcal{X}\supseteq L^{\infty}$. We find an expression for the inf-convolution and also the corresponding optimal allocation, covering the results of the inf-convolution of $\Lambda\mathrm{VaR}$ or $\mathrm{VaR}$ with homogeneous or heterogenous beliefs in Liu (2024), Embrechts et al. (2020) and Embrechts et al. (2018). We first consider the general beliefs and then discuss three different cases about the relation of the beliefs: homogeneous beliefs, conditional beliefs and absolutely continuous beliefs.

### v3

@@ Abstract
In this paper, we study the risk sharing problem among multiple agents using Lambda Value-at-Risk as their preference functional, under heterogeneous beliefs, where beliefs are represented by several probability measures. We obtain semi-explicit formulas for the inf-convolution of multiple Lambda Value-at-Risk measures under heterogeneous beliefs and the explicit forms of the corresponding optimal allocations. To show the impact of belief heterogeneity, we consider three cases: homogeneous beliefs, conditional beliefs, and general beliefs with two agents. For those cases, we find more explicit expressions for the inf-convolution, showing the influence of the relation of the beliefs on the inf-convolution. Moreover, we consider, in a two-agent setting, the inf-convolution of one Lambda Value-at-Risk and a general risk measure, including expected utility, distortion risk measures and Lambda Value-at-Risk as special cases, with differing beliefs. The expression of the inf-convolution and the form of the optimal allocation are obtained. In all above cases we demonstrate that trivial outcomes arise when both belief inconsistency and risk tolerance are high. Finally, we discuss risk sharing for an alternative definition of Lambda Value-at-Risk.
@@ 1 Introduction
The relation between the beliefs plays a crucial role in the risk sharing problem; see e.g., Chi (2019). To study the impact of belief heterogeneity, we consider three cases in Section 3: homogeneous beliefs, conditional beliefs, and general beliefs with two agents. For homogeneous beliefs, we show that the inf-convolution of $\Lambda\mathrm{VaR}$ is still a $\Lambda\mathrm{VaR}$ for general $\Lambda$ functions, extending the results in Liu (2025) and Xia and Hu (2024), where the monotonicity of the $\Lambda$ functions is typically required. In particular, our results include the case that some $\Lambda\mathrm{VaR}$ have increasing $\Lambda$ functions and others have decreasing $\Lambda$ functions, demonstrating that the agents may have relatively different risk appetites. The conditional beliefs reflect the asymmetric information obtained by the agents such that each agent has slightly different concern. Under this setup, we also obtain an explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$, showing that the inf-convolution of $\Lambda\mathrm{VaR}$ under conditional beliefs is also a $\Lambda\mathrm{VaR}$ under a new conditional belief. Conditional beliefs are closely related to the conditional risk measures proposed in the literature to quantify systemic risk, such as $\mathrm{CoVaR}$ (Adrian and Brunnermeier (2016) and Girardi and Tolga Ergün (2013)) and $\mathrm{CoES}$ (Mainik and Schaanning (2014)). The risk sharing problem in this case can be interpreted as risk sharing with some Co-risk measures ($\mathrm{Co\Lambda\mathrm{VaR}}$, a new conditional risk measure proposed in Section 3). For the third case, we consider the risk sharing between two agents. With the aid of the Lebesgue decomposition theorem, we find a more explicit formula for the inf-convolution of $\Lambda\mathrm{VaR}$ with general heterogeneous beliefs. Note that we also give the explicit expression for the inf-convolution of $\mathrm{VaR}$ with heterogeneous beliefs, which are also new to the literature.
[…]
@@
The notation and key definitions are given in Section 2 and all the proofs are delegated to Appendix A.
@@ 2 Notation and Definitions
For a given probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $L^{\infty}$ denote the collection of all bounded random variables. Moreover, let $\mathcal{X}$ be a set of random variables containing $L^{\infty}$. We say $\mathcal{X}$ is unbounded if $\mathcal{X}$ contains unbounded random variables. We suppose $\mathcal{X}$ has good enough properties to conduct our study, such as $\mathcal{X}=L^{p}$ for some $p\geqslant 0$, where $L^{p}$ represents the set of all random variables with finite $p$-th moments. Throughout this paper, we consider the random variables themselves but not the corresponding equivalent classes, because, as will be seen in the sequel, more than one probability measures are considered simultaneously.
For any $X\in\mathcal{X}$, a positive value of $X$ represents a financial loss. For a risk measure, that is, a mapping $\rho:\mathcal{X}\to\mathbb{R}$, we say that $\rho$ monotone if $X\leqslant Y$ implies $\rho(X)\leqslant\rho(Y)$; and $\rho$ is cash-additive if $\rho(X+c)=\rho(X)+c$ for $X\in\mathcal{X}$ and $c\in\mathbb{R}$. We say $\rho$ is a monetary risk measure if it is monotone and cash-additive. Furthermore, $\rho$ is law-invariant under a probability measure $\mathbb{Q}$ if for all $X,Y\in\mathcal{X}$,
@@
[…]
For a probability measure $\mathbb{Q}$, the distribution function of $X$ under $\mathbb{Q}$ is denoted as $F_{X}^{\mathbb{Q}}$. For $X\in\mathcal{X}$, $F_{X}^{\mathbb{Q},-1}$ represents its left-quantile under $\mathbb{Q}$, which is defined by
@@
with the convention that $\inf\emptyset=\infty$. For any $X\in\mathcal{X}$, if $\mathbb{Q}$ is atomless, we denote by $U_{X}^{\mathbb{Q}}$ a uniform random variable on $[0,1]$ under $\mathbb{Q}$ such that $X=F_{X}^{\mathbb{Q},-1}(U_{X}^{\mathbb{Q}})$ a.s. under $\mathbb{Q}$. The existence of such $U_{X}^{\mathbb{Q}}$ for any random variable $X$ is guaranteed by e.g., Lemma A.32 of Föllmer and Schied (2016).
We next define tail risk measures, which are important for our results later. For a random variable $X\in\mathcal{X}$, $p\in(0,1]$ and an atomless probability measure $\mathbb{Q}$, we call
$X_{p}=(F_{X}^{\mathbb{Q}})^{-1}(1-p+pU_{X}^{\mathbb{Q}})$
the tail risk of $X$ beyond its (1p)(1-p)-quantile under $\mathbb{Q}$. The distribution of $X_{p}$ under $\mathbb{Q}$ is given by
$\mathbb{Q}(X_{p}\leqslant x)=\frac{(F_{X}^{\mathbb{Q}}(x)-(1-p))_{+}}{p},~x\in\mathbb{R},$
where $x_{+}=\max(x,0)$. We say $\rho$ is a $p$-tail risk measure under $\mathbb{Q}$ for some $p\in(0,1)$ if $X_{p}\overset{\mathbb{Q}}{=}Y_{p}$ implies $\rho^{\mathbb{Q}}(X)=\rho^{\mathbb{Q}}(Y)$ for all $X,Y\in\mathcal{X}$. One can refer to Liu and Wang (2021) and Liu et al. (2022) for more details on the definition, properties and applications of tail risk measures.
@@
Throughout this paper, we assume that the probability measures $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ live on $(\Omega,\mathcal{F})$ and, with the exception of Subsection 3.1, the corresponding probability spaces are atomless. In this paper, we suppose the agents may have heterogeneous beliefs, represented by some probability measures $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$. The inf-convolution of the risk measures $\rho_{1},\dots,\rho_{n}$, under the respective probability measures $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ is the mapping $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}:\mathcal{X}\to[-\infty,\infty)$, defined as
@@
Note that $\rho_{i}$ is law-invariant under $\mathbb{Q}_{i}$, where $\mathbb{Q}_{i}$ represents the belief of the i-th agent. In case of belief heterogeneity, the inf-convolution may not be law-invariant, i.e., if $X\overset{\mathbb{Q}_{i}}{=}Y$ for all $i=1,\dots,n$, then $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)\neq\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(Y)$ may hold; see more details in the Remark at the end of Section 3.3.
An $n$-tuple $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is called an optimal allocation of $X$ for $(\rho_{1},\dots,\rho_{n})$ under $(\mathbb{Q}_{1},\dots,\mathbb{Q}_{n})$ if $\sum_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X_{i})=\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)$. The inf-convolution $\mathop{\square}\displaylimits_{i=1}^{n}\rho_{i}^{\mathbb{Q}_{i}}(X)$ can be interpreted as the smallest possible aggregate capital for the total risk $X$ in financial system, if $\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ represents the capital charge for the i-th financial institution to hold the risky position $X_{i}$. More economic interpretations on the inf-convolution can be found in e.g., Delbaen (2012), Rüschendorf (2013) and Embrechts et al. (2018).
For the risk measures $(\rho_{1},\dots,\rho_{n})$ under $(\mathbb{Q}_{1},\dots,\mathbb{Q}_{n})$ and a total risk $X$, an allocation $(X_{1},\dots,X_{n})\in\mathbb{A}_{n}(X)$ is Pareto-optimal if for any other allocation $(Y_{1},\dots,Y_{n})\in\mathbb{A}_{n}(X)$, $\rho_{i}^{\mathbb{Q}_{i}}(Y_{i})\leqslant\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ for all $i=1,\dots,n$ implies $\rho_{i}^{\mathbb{Q}_{i}}(Y_{i})=\rho_{i}^{\mathbb{Q}_{i}}(X_{i})$ for all $i=1,\dots,n$. For finite-valued monetary risk measures, it is shown in Embrechts et al. (2018) that an allocation is optimal if and only if it is Pareto-optimal. Note that $\Lambda\mathrm{VaR}$ (defined in (2)) does not satisfy cash-additivity; see Frittelli et al. (2014) and Bellini and Peri (2022). Hence it is not a monetary risk measure. However, one can still show that the optimal allocation of the inf-convolution of $\Lambda\mathrm{VaR}$ is Pareto-optimal.
@@
where $\inf\emptyset=\infty$ and $\sup\emptyset=$-$\infty$. Note that $\Lambda\mathrm{VaR}^{\mathbb{Q}}(X)=\Lambda\mathrm{VaR}^{+,\mathbb{Q}}(X)$ if $\Lambda$ is increasing; Otherwise, it may not be true; see Proposition 6 of Bellini and Peri (2022). We refer to Bellini and Peri (2022) for two other definitions of $\Lambda\mathrm{VaR}$. If $\Lambda$ is a constant, then $\Lambda\mathrm{VaR}^{\mathbb{Q}}$ boils down to $\mathrm{VaR}^{\mathbb{Q}}$, i.e., $\mathrm{VaR}^{\mathbb{Q}}$ at level $p\in[0,1)$ is given by
@@
The conclusion in Proposition 1 implies that for any two monetary risk measures $\rho_{1}$ and $\rho_{2}$, $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)=$-$\infty$ if $\mathbb{Q}_{1}$ and $\mathbb{Q}_{2}$ are mutually singular. This means that if the beliefs of agents are too divergent, then each one of them can get unreasonably good deal by transferring risks between them. Then, the fact that the inf-convolution is -$\infty$ reflects that agents can always make their position better and hence there is no optimal allocation. In contrast, if $\mathbb{Q}_{1}=\mathbb{Q}_{2}$, then $\rho_{1}^{\mathbb{Q}_{1}}\mathop{\square}\displaylimits\rho_{2}^{\mathbb{Q}_{2}}(X)>$-$\infty$ for many monetary risk measures $\rho_{1}$ and $\rho_{2}$; see e.g., Embrechts et al. (2018) for quantile-based risk measures and Filipović and Svindland (2008) for convex risk measures. Note that the conclusion in Proposition 1 can be easily extended to the case with $n$ agents for $n\geqslant 3$.
@@ 3 Inf-convolution of multiple $\Lambda\mathrm{VaR}$
In this section, we investigate the inf-convolution of $n$ Lambda Value-at-Risk with heterogeneous beliefs represented by $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$, respectively. Suppose $\mathcal{X}\supseteq L^{\infty}$. We find an expression for the inf-convolution and also the corresponding optimal allocation, covering the results of the inf-convolution of $\Lambda\mathrm{VaR}$ or $\mathrm{VaR}$ with homogeneous or heterogeneous beliefs in Liu (2025), Embrechts et al. (2020) and Embrechts et al. (2018). We first consider the general beliefs and then discuss three different cases about the relation of the beliefs: homogeneous beliefs, conditional beliefs, and general beliefs with two agents.